# BST: Badminton Stroke-type Transformer for Skeleton-based Action Recognition in Racket Sports

### v1

@@ Abstract
Badminton, known for having the fastest ball speeds among all sports, presents significant challenges to the field of computer vision, including player identification, court line detection, shuttlecock trajectory tracking, and player stroke-type classification. In this paper, we introduce a novel video segmentation strategy to extract frames of each player’s racket swing in a badminton broadcast match. These segmented frames are then processed by two existing models: one for Human Pose Estimation to obtain player skeletal joints, and the other for shuttlecock trajectory detection to extract shuttlecock trajectories. Leveraging these joints, trajectories, and player positions as inputs, we propose Badminton Stroke-type Transformer (BST) to classify player stroke-types in singles. To the best of our knowledge, experimental results demonstrate that our method outperforms the previous state-of-the-art on the largest publicly available badminton video dataset, ShuttleSet, which shows that effectively leveraging ball trajectory is likely to be a trend for racket sports action recognition.
@@ 2.2 Skeleton-based Action Recognition (SAR)
SkateFormer [9], proposed by Jeonghyeok Do and Munchurl Kim, builds upon the Transformer architecture with several modifications. They integrated GCN, TCN, and their custom-designed Skate-MSA into the self-attention module. The core idea of Skate-MSA is to transform the input data into four different types, which were constructed based on their observations of human motion patterns. Each type is then processed separately. Their work was later published at ECCV 2024.
Hongda Liu et al. proposed ProtoGCN [22], which is also based on a custom GCN and a multi-scale TCN. Additionally, they incorporated Class-Specific Contrastive Loss to enhance model training. Their approach achieved strong performance on several well-known public datasets in SAR.
@@ 3.1 Video Segmentation Strategy
for $i=1,2,\ldots,r$, where $r$ is the total number of rallies in the match video, and $j=1,2,\ldots,n^{(i)}$. We now segment every rally into stroke clips. Suppose a match video contain $N$ frames denoted by $F_{1},F_{2},\ldots,F_{N}$. Some of the frames in the match video are hit frames, which are the frames in which the shuttlecock is in contact with, or closest to, the racket during a stroke. Thus, every stroke $S^{(i)}_{j}$ in rally $R_{i}$ has a one-to-one correspondence with the hit frame number $h^{(i)}_{j}$. A fixed-width clipping strategy similar to the approach used in [35] define the $j$-th stroke clip in rally $R_{i}$ to be
@@
for $j=1,2,\ldots,n^{(i)}$ and $1\leq i\leq r$. In Eq. 1, $t$ is ideally half of the width of every hit frame number. However, it is a fixed parameter, and clearly, all stroke clips defined in Eq. 1 realistically contain at least one hit frame and possibly many hit frames.
@@
If the target stroke being classified is the first or last stroke in a rally, we set the lower or upper bound of the stroke clip to be the same value as used in Eq. 1. In Eq. 2, $\epsilon$ is a small parameter less than $t$. Looking at this strategy from the perspective of shuttlecock trajectory, it captures clean and complete trajectories in three stages: (1) the shuttlecock flying towards the target player’s court, (2) the shuttlecock trajectory after being hit by the target player, and (3) the direction of the shuttlecock trajectory after the opponent’s response. These stages respectively represent (1) the target player’s observation and reaction phase, (2) the actual trajectory of the target stroke, and (3) the opponent’s response pattern. This allows the model to not only judge based on (2) but also understand the target player’s common strategies after going through (1) and infer the target stroke-type back from (3). The reason that the stroke information in (3) is less than that in (1), which means that $\epsilon$ cannot be set too large, is to preserve the characteristic of keeping the target hit frame close to the middle of the stroke clip under the ideal condition in Eq. 1.
We remark that the dataset (Sec. 4.1) we used to train our model contains hit frame and rally information. To classify a match video without hit frame and rally information, we can use a model called HitNet proposed in MonoTrack [23] to detect hit frames. To detect rallies, we note that time intervals between successive rallies are in general much larger than the time differences between two typical successive hit frames in one rally. In practice, we also set some limits on $\hat{h}^{(i)}_{j-1}$ and $\hat{h}^{(i)}_{j+1}$ to prevent the model from accessing irrelevant information caused by remaining unclean data.
@@ 3.3.1 BST-0
where $X^{(l)}\in\mathbb{R}^{S\times D_{model}}$ is the output to the $l$-th layer, $S$ is the sequence length, $D_{model}$ is the dimension used in the model, LN denotes the layer normalization, MHSA is the multi-head self-attention mechanism, and FFN is the feedforward network. Before passing through the MHSA, the input $X^{(l)}$ which simplifies here to $X$ is splitted into:
@@
Next, a key concept of our approach is ensuring that the model focuses on extracting critical information from the shuttlecock trajectory. To achieve this, we design a Cross Transformer layer with a Multi-Head Cross Attention (MHCA), which differs from MHSA in one key aspect: in the attention mechanism, $K$ and $V$ are derived from the shuttlecock trajectory latent, while $Q$ is generated from the player pose latent or the player position latent, respectively, to perform cross-attention. As a result, the subsequent outputs follow two distinct pathways. One pathway is dedicated to the trajectory information related to the first player poses, while the other pathway focuses on the trajectory information related to the second player poses. Both pathways are fed into a second Transformer Encoder.
Finally, the model generates logits through an MLP (multi-layer perceptron) head, which are converted into probability values using the softmax function. The model is trained using cross-entropy loss.
@@ 3.3.4 BST-2
In BST-2, shown in Fig. 5, we compare the cosine similarity between input (1) and (2) with that between input (1) and (3). The difference between these two values gives $\tilde{\alpha}$. Since cosine similarity ranges from -1 to 1, $\tilde{\alpha}\in[-2,2]$. We normalize it using $\alpha=(\tilde{\alpha}+2)/4$ so that $\alpha\in[0,1]$. $\alpha$ denotes the coefficient on the target player information, so $1-\alpha$ is the coefficient on his/her opponent’s information. This ensures that the player with a stronger correlation to the overall shuttlecock trajectory exerts a greater influence on the input of the MLP head.
@@ 4.1 Dataset
The original ShuttleSet dataset consists of 19 distinct stroke categories, including a ”None” type. One category, driven flight, containing fewer than 50 strokes across the entire dataset was merged into the ”None” category by us. Additionally, since the dataset does not differentiate between strokes played by the top or bottom player within each category (except for ”None”), the number of stroke categories (except for ”None”) was doubled. This adjustment resulted in a total of 35 categories, including ”None”.
4.2 Results
@@
Finally, the incorporation of our segmentation strategy results in even greater performance gains across all BST models, significantly exceeding the performance of TemPose, as shown in Tab. 2.

### v2

@@ Abstract
Badminton, known for having the fastest ball speeds among all sports, presents significant challenges to the field of computer vision, including player identification, court line detection, shuttlecock trajectory tracking, and player stroke-type classification. In this paper, we introduce a novel video segmentation strategy to extract frames of each player’s racket swing in a badminton broadcast match. These segmented frames are then processed by two existing models: one for Human Pose Estimation to obtain player skeletal joints, and the other for shuttlecock trajectory detection to extract shuttlecock trajectories. Leveraging these joints, trajectories, and player positions as inputs, we propose Badminton Stroke-type Transformer (BST) to classify player stroke-types in singles. To the best of our knowledge, experimental results demonstrate that our method outperforms the previous state-of-the-art on the largest publicly available badminton video dataset, ShuttleSet, which shows that effectively leveraging ball trajectory is likely to be a trend for racket sports action recognition. The code is publicly available at https://github.com/Va6lue/BST-Badminton-Stroke-type-Transformer.
@@ 2.2 Skeleton-based Action Recognition (SAR)
SkateFormer [9], proposed by Jeonghyeok Do and Munchurl Kim, builds upon the Transformer architecture with several modifications. They integrated GCN, TCN, and their custom-designed Skate-MSA into the self-attention module. The core idea of Skate-MSA is to transform the input data into four different types, which were constructed based on their observations of human motion patterns. Each type is then processed separately. Their work was published at ECCV 2024.
Hongda Liu et al. proposed ProtoGCN [22], which is also based on a custom GCN and a multi-scale TCN. Additionally, they incorporated Class-Specific Contrastive Loss to enhance model training. Their approach achieved strong performance on several well-known public datasets in SAR and was later published at CVPR 2025.
@@ 3.1 Video Segmentation Strategy
for $i=1,2,\ldots,r$, where $r$ is the total number of rallies in the match video, and $j=1,2,\ldots,n^{(i)}$. We now segment every rally into stroke clips. Suppose a match video contains $N$ frames denoted by $F_{1},F_{2},\ldots,F_{N}$. Some of the frames in the match video are hit frames, which are the frames in which the shuttlecock is in contact with, or closest to, the racket during a stroke. Thus, every stroke $S^{(i)}_{j}$ in rally $R_{i}$ has a one-to-one correspondence with the hit frame number $h^{(i)}_{j}$. A fixed-width clipping strategy similar to the approach used in [35] define the $j$-th stroke clip in rally $R_{i}$ to be
@@
for $j=1,2,\ldots,n^{(i)}$ and $1\leq i\leq r$. In Eq. 1, from the perspective of having comprehensive player pose information, $t$ ensures that the target hit frame $F_{h^{(i)}_{j}}$ is at the middle of the stroke clip $C^{(i)}_{j}$ and is ideally set to half of the interval between every two consecutive hit frame numbers. However, it is a fixed parameter, and clearly, all stroke clips defined in Eq. 1 realistically contain at least one hit frame and possibly many hit frames.
@@
If the target stroke being classified is the first or last stroke in a rally, we set the lower or upper bound of the stroke clip to be the same value as used in Eq. 1. In Eq. 2, $\epsilon$ is a small parameter less than $t$. Looking at this strategy from the perspective of shuttlecock trajectory, it captures clean and complete trajectories in three stages: (1) the shuttlecock flying towards the target player’s court, (2) the shuttlecock trajectory after being hit by the target player, and (3) the direction of the shuttlecock trajectory after the opponent’s response. These stages respectively represent (1) the target player’s observation and reaction phase, (2) the actual trajectory of the target stroke, and (3) the opponent’s response pattern. This allows the model to not only judge based on (2) but also understand the target player’s common strategies after going through (1) and infer the target stroke-type back from (3). Thus, our goal is achieved, and the most informative segment of the shuttlecock trajectory related to the target stroke is captured. The reason that the stroke information in (3) is less than that in (1), which means that $\epsilon$ cannot be set too large, is to preserve the characteristic of keeping the target hit frame close to the middle of the stroke clip under the ideal condition in Eq. 1.
[…]
In practice, we also set some limits on $\hat{h}^{(i)}_{j-1}$ and $\hat{h}^{(i)}_{j+1}$ to prevent the model from accessing irrelevant information caused by remaining unclean data. Specifically, the actual time distance of these two frame numbers from the target frame number does not exceed 1.5 seconds (before adding $\epsilon$). Additionally, $t$ is set to the half of the FPS of the video representing 0.5 seconds interval, and $\epsilon$ is set to $\frac{t}{2}$.
@@ 3.3.1 BST-0
where $X^{(l)}\in\mathbb{R}^{S\times D_{model}}$ is the output to the $l$-th layer, $S$ is the sequence length, $D_{model}$ is the dimension used in the model, LN denotes the layer normalization, MHSA is the multi-head self-attention mechanism, and FFN is the feedforward network. Before passing through the MHSA, the input $X^{(l-1)}$ which simplifies here to $X$ is split into:
@@
Next, a key concept of our approach is ensuring that the model focuses on extracting critical information from the shuttlecock trajectory. To achieve this, we design a Cross Transformer layer with a Multi-Head Cross Attention (MHCA), which differs from MHSA in one key aspect: in the attention mechanism, $K$ and $V$ are derived from the shuttlecock trajectory latent, while $Q$ is generated from the player pose latent or the player position latent, respectively, to perform cross-attention. As a result, the subsequent outputs follow two distinct pathways. One pathway is dedicated to the trajectory information related to the first player poses, while the other pathway focuses on the trajectory information related to the second player poses. Both pathways are fed into a second Transformer Encoder. Finally, the model generates logits through an MLP (multi-layer perceptron) head, which are converted into probability values using the softmax function.
Trans1 and Trans2 are inspired by TemPose [16] (Temporal and Interactional Transformer), but Trans2 is actually different from their Interactional Transformer that serves people as a sequence. Our Trans2 still processes the frame sequence to obtain the class token containing the trajectory information related to player 1 or 2. For CrossTrans, our idea involves calculating the similarity between pose and trajectory, and using these similarity values to pick the value arrays in $V$ (from trajectory), which is a critical information from our perspective, an insight gained after observing our experimental results in Sec. 4 (TemPose-PF vs. TemPose-SF in Tab. 1).
In practice, there are 2 layers in Trans1, 1 layer in CrossTrans, and 1 layer in Trans2.
@@ 3.3.4 BST-2
In BST-2, shown in Fig. 5, we compare the cosine similarity between input (1) and (2) with that between input (1) and (3). The difference between these two values gives $\tilde{\alpha}$. Since cosine similarity ranges from -1 to 1, $\tilde{\alpha}\in[-2,2]$. We normalize it using $\alpha=(\tilde{\alpha}+2)/4$ so that $\alpha\in[0,1]$. $\alpha$ denotes the coefficient on the player one’s information, so $1-\alpha$ is the coefficient on the player two’s information. This ensures that the player with a stronger correlation to the overall shuttlecock trajectory exerts a greater influence on the input of the MLP head.
@@ 4 Experiments
Model Modality SP PP Acc Macro-F1 Acc-2
TemPose-SF J+B × 0.591 0.453 0.781
BST-0 J+B × 0.642 0.512 0.812
TemPose-TF [16] J+B 0.601 0.445 0.775
BST-1 J+B 0.640 0.538 0.828
BST-2 J+B 0.636 0.518 0.820
BST-3 J+B 0.640 0.534 0.825
Table 3: Model comparison trained on 25% of the training set without our segmentation strategy.
@@ 4.1 Dataset
The original ShuttleSet dataset consists of 19 distinct stroke categories, including a ”none” type. One category, driven flight, containing fewer than 50 strokes across the entire dataset was merged into the ”none” category by us. Additionally, since the dataset does not differentiate between strokes played by the top or bottom player within each category (except for ”none”), the number of stroke categories (except for ”none”) was doubled. This adjustment resulted in a total of 35 categories, including ”none”. (See Tab. A for the complete list of classes.)
4.2 Main Results
@@
Finally, the incorporation of our segmentation strategy results in even greater performance gains on the metrics across BST models, significantly exceeding the performance of TemPose using our segmentation strategy, as shown in Tab. 2.
For a deeper look into the results between Tab. 1 and Tab. 2, we can observe two phenomena. First, the performance of TemPose drops when using our segmentation strategy. Our segmentation strategy creates variable sequence lengths due to the fluctuating match pace. TemPose struggles with these, possibly due to limitations in its Interactional Transformer, which is designed to process people as a sequence. (We have mentioned this in the last paragraph of Sec. 3.3.1.) The other phenomenon is that when comparing the results of BSTs themselves, the differences between BSTs (without vs. with our segmentation strategy) from BST-1 to BST-3 are [0.23%, 0.22%, 1.03%], [-0.02%, -0.23%, 0.52%], [0.15%, -0.01%, 0.73%]. Notably, 2 of the 3 observed performance drops occur with BST-2. This is because BST-2’s pathways, prior to the MLP head, are limited to two branches (trajectory information from player 1 and from player 2), lacking the overall trajectory information. Furthermore, our segmentation strategy primarily focuses on shuttlecock trajectory information. Consequently, the performance drop is more pronounced for BST-2 than for BST-3. Despite a modest drop on 3 out of 9 metrics, BSTs still benefit from our segmentation strategy, especially with high class granularity (See Sec. 5.1 for more details), where Acc-2 becomes more important. Moreover, compared to the best of TemPose in Tab. 1, BSTs outperform it on 7 out of 9 metrics, with one loss being only -0.01%. Lastly, our prototype BST-0, even without PP input but utilizing our segmentation strategy, still beats the best of TemPose on Acc-2 (0.9306 vs. 0.9242).
4.3 Results on Reduced Training Set
Tab. 3 shows the results of models trained on only (each class) 25% of the training set and tested on the same testing set. This indicates that our BST architecture demonstrates superior generalization capability compared to TemPose [16] when the training dataset is small and reference information is limited.
@@ 6 Conclusion
Supplementary Material
A Training Speed
On a NVIDIA RTX 4070 Ti Super setup, training BST-3 takes 29 sec per epoch, while training TemPose-TF takes 31 sec per epoch. Fig. A also shows that training BST-3 converges faster than TemPose-TF.
Figure A: Loss Curves of BST-3 and TemPose-TF.
B Training Settings
The implementation is based on the PyTorch framework. All downstream classification models in Tab. 1 were trained from scratch using cross-entropy loss. Moreover, the compared models are derived from their official public implementations for better code readability and more efficient training and inference. We chose AdamW with default parameters as the optimizer and cosine annealing with warm-up learning rate scheduler to adjust the learning rate during training. The hyperparameters used in our experiments are listed below:
  • n_epochs: $1600$
  • early_stop_n_epochs: $300$
  • batch_size: $128$
  • learning_rate: $5\times 10^{-4}$
  • warm_up_step: $400$
  • cosine_annealing_num_cycles: $0.25$
  • n_classes: $35$
[…]
C Classes in the Dataset
There are $35$ classes in the dataset as shown in Tab. A.
Class ID Num of Strokes Player 中文類別名稱 Class Name
0 2982 Top 放小球 net shot
1 1618 Top 擋小球 return net
2 1287 Top 殺球 smash
3 831 Top 點扣 wrist smash
4 2343 Top 挑球 lob
5 123 Top 防守回挑 defensive return lob
6 1309 Top 長球 clear
7 319 Top 平球 drive
8 234 Top 後場抽平球 back-court drive
9 1036 Top 切球 drop
10 615 Top 過渡切球 passive drop
11 1317 Top 推球 push
12 270 Top 撲球 rush
13 144 Top 防守回抽 defensive return drive
14 606 Top 勾球 cross-court net shot
15 904 Top 發短球 short service
16 159 Top 發長球 long service
17 2842 Bottom 放小球 net shot
18 1756 Bottom 擋小球 return net
19 1075 Bottom 殺球 smash
20 728 Bottom 點扣 wrist smash
21 2536 Bottom 挑球 lob
22 155 Bottom 防守回挑 defensive return lob
23 1352 Bottom 長球 clear
24 335 Bottom 平球 drive
25 201 Bottom 後場抽平球 back-court drive
26 943 Bottom 切球 drop
27 583 Bottom 過渡切球 passive drop
28 1335 Bottom 推球 push
29 201 Bottom 撲球 rush
30 238 Bottom 防守回抽 defensive return drive
31 620 Bottom 勾球 cross-court net shot
32 954 Bottom 發短球 short service
33 200 Bottom 發長球 long service
34 1278 - 未知球種 none
Table A: Classes in the dataset.